Madison and Alexa have a combined age of 32. Madison is older, and the difference in
their ages is 4 years. How old is each person?

Found 2 solutions by Boreal, Edwin McCravy:
Answer by Boreal(15235)

A=x
M=x+4 so difference is 4
they add (2x+4) to 32 (=32)
2x=28
x=14 A
x+4=18 M
(units years)

Answer by Edwin McCravy(20059)


 M + A = 32
 M - A =  4
-----------
2M     = 36     <--adding the two equations term by term
 M     = 18     <--dividing both sides by 2

 M + A = 32     
18 + A = 32     <--substituting 18 for M, from above
     A = 14     <--subtracting 18 from both sides 

Edwin